OpenStudy (anonymous):

What is the volume of a square pyramid with base edges of 20 meters and a height of 30 meters? 600 m3 12,000 m3 4,000 m3 1,800 m3

OpenStudy (anonymous):

volume = 1/3 base^2 h. volume = 1/3 20^2 * 30 ~

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@nphuongsun93 ?

OpenStudy (anonymous):

I mean C ?

ganeshie8 (ganeshie8):

its B

OpenStudy (anonymous):

how ? 1/3 * 400* 30 is 4,000 ?

ganeshie8 (ganeshie8):

huh i did wrong arithmetic :( you're right :) its C

OpenStudy (anonymous):

:D